# Modelling the 5G Energy Consumption using Real-world Data: Energy Fingerprint is All You Need

### v1

@@ Abstract
The introduction of fifth-generation (5G) radio technology has revolutionized communications, bringing unprecedented automation, capacity, connectivity, and ultra-fast, reliable communications. However, this technological leap comes with a substantial increase in energy consumption, presenting a significant challenge. To improve the energy efficiency of 5G networks, it is imperative to develop sophisticated models that accurately reflect the influence of base station (BS) attributes and operational conditions on energy usage. Importantly, addressing the complexity and interdependencies of these diverse features is particularly challenging, both in terms of data processing and model architecture design.
This paper proposes a novel 5G base stations energy consumption modelling method by learning from a real-world dataset used in the ITU 5G Base Station Energy Consumption Modelling Challenge in which our model ranked second. Unlike existing methods that omit the Base Station Identifier (BSID) information and thus fail to capture the unique energy fingerprint in different base stations, we incorporate the BSID into the input features and encoding it with an embedding layer for precise representation. Additionally, we introduce a novel masked training method alongside an attention mechanism to further boost the model’s generalization capabilities and accuracy. After evaluation, our method demonstrates significant improvements over existing models, reducing Mean Absolute Percentage Error (MAPE) from 12.75% to 4.98%, leading to a performance gain of more than 60
@@ I Introduction
The advent of the fifth-generation (5G) network introduces advanced capabilities that significantly benefit various applications, including virtual reality, vehicular networks and the Internet of Things (IoT). However, this technological leap comes with a substantial increase in energy consumption. Compared to its predecessor, the fourth-generation (4G) network, the energy consumption of the 5G network is approximately three times higher [1]. Notably, energy costs constitute about 22.5% of the total expenditure for mobile network operators [2], a figure that is on the rise. A critical factor in this context is the radio access network (RAN), particularly the base stations (BSs), which account for over 70% of the network’s energy usage [3].
Commercial networks feature BSs with various architectures, such as remote radio units or antenna processing units. The power consumption profile of these BSs, which are equipped with diverse capabilities such as varying numbers of carriers and antennas, are further influenced by their configuration settings, impacting their energy consumption. Despite their substantial energy requirements, there exist methods to mitigate RAN consumption, primarily by disabling specific hardware components within the BSs. These energy-saving methods can be activated manually or automatically under certain traffic conditions. Understanding the interplay of these factors is essential for optimizing future RAN energy efficiency.
The precise assessment of BS energy consumption is pivotal in the modelling of energy consumption in 5G network, a topic that has been a longstanding challenge and has consistently garnered attention, with numerous studies conducted to investigate this issue. The work presented in [4] introduced a popular analytical power consumption model highlighting the linear correlation between the power output of the BS and its power usage. This model is further extended in [5], incorporating considerations for the massive multiple-input multiple-output (mMIMO) architecture and multiple energy-efficient techniques. Subsequent work in [6] expanded the model in [4] considering a linear increase of the power consumption with the number of mMIMO transceivers. A comprehensive and detailed framework for understanding power consumption was proposed in [7], taking into account factors like the mMIMO architecture, downlink and uplink phases, and the number of users multiplexed.
@@
So far, several emerging tasks such as sensing [8], mobile edge computing [9], edge learning and inference [10, 11] have been applied to extend BS applications for next-generation networks. These additional tasks not only increase the BSs’ energy consumption but also introduce new characteristics relevant to energy consumption modelling, directly impacting how energy efficiency is understood and optimized. The increasing complexity of mobile networks, coupled with the diverse capabilities of BSs, necessitates the use of machine learning models that can effectively account for the impact of various parameters on energy consumption. Most recently, SRCON (Simulated Reality of Communication Networks) has emerged as a revolutionary paradigm in network simulation. Detailed in  [12, 13], SRCON leverages an innovative mix of white-box and black-box models to mimic the stochastic behaviors observed in real-world 4G/5G mobile networks, offering a robust framework for accurate simulation. Within this framework, [14, 15] develops a machine learning model, trained on extensive network measurement data, which significantly improves the accuracy of energy consumption estimations compared to traditional mathematical modelling.
However, existing modelling methods do not take the complexity and heterogeneity of real commercial products into consideration. Such complexity is reflected in various previously unconsidered features, such as hardware characteristics and physical environments, which we refer to as the energy fingerprint. Omitting the said energy fingerprint may lead to the incomplete input feature space and, what’s worse, the learnt model may suffer from severe one-to-many issue, i.e., the same input features may be mapped to different prediction outcomes, posing significant challenges to achieving accurate prediction outcomes.
To alleviate the one-to-many issue, we propose a novel modelling method based on the real-world dataset from the ITU 5G Base Station Energy Consumption Modelling Challenge [16] where our solution ranked second among more than two hundred teams throughout the world. Specifically, our model incorporates the Base Station Identifier (BSID) feature as input, which was ignored by previous methods, to capture the inherent energy fingerprint of different BSs. Furthermore, we utilize an embedding layer for BSID feature. This method allows a more precise representation of the BSID feature and addresses the shortcomings of the classic One-Hot encoding, which leads to a prohibitively high-dimension of the input feature space and thus an unnecessarily large model size of the neural network to accommodate the high-dimension input, hindering the model’s ability to generalize accurately. More importantly, it is found that the embedding layer can project the BSID into a latent space in which BSs are clustered according to their energy fingerprint, justifying the capability of the proposed model in energy fingerprint extraction. Moreover, to endow the embedding layer with generalization capability to handle the unseen BSID in the test data, we further introduce a new masked training technique to boost the model’s generalization ability, making it more applicable in real-world settings. Last but not least, we design the attention mechanism for further automatic feature selection besides the manual selection before training. Extensive experiments on real-world 5G network measurement data are conducted to examine the superiority of the proposed design over the existing baselines in diverse test scenarios. The Mean Absolute Percentage Error (MAPE) improved from 12.75% in the existing method to 4.98%, resulting in a performance enhancement of over 60% for the algorithm.
@@
In this section, we introduce the 5G Radio Unit Dataset. The dataset used in our study comprised 102,705 hourly measurements collected over a period of 8 days. Each sample encapsulates a variety of features, classified into four main categories:
@@
The dataset is split into training and testing sets to test our model’s generalization capabilities, with certain BSs only present in the testing set. The training set consists of 92,629 samples. The testing set includes 10,076 samples, of which 3,067 are from BSs not encountered during training (unknown BS), and 7,009 are from BSs that appear in the training set (known BS). All BS refers to the entirety of both known and unknown BSs combined. The training data includes 10 distinct RUType categories, while the testing set introduces two additional RUTypes not seen in the training, challenging the model’s ability to adapt to new RUTypes.
@@ III Problem Statement
By analysis of the collected data, there exists a “one-to-many” issue encountered in the datasets, meaning that different energy consumption values are reported for identical input feature values. To better illustrate the said issue, we depict Fig. 1, a scatter plot, in which the labels (i.e., the energy consumption values) of selected data samples are plotted against one of the most important features (i.e., the load of the primary cell as identified in [15]) with other features kept identical.
Three key observations can be made from Fig. 1 as listed below:
  • Energy consumption exhibits a roughly linear relationship with the load of the primary cell.
  • Without distinguishing BSID, as depicted by the gray area in the figure, a common load value corresponds to high variance in energy consumption.
@@
Combining these observations, we can conclude that distinguishing BSID helps a lot in reducing the variance of energy consumption predictions. This gain comes from the fact that incorporating BSID as model input feature enables us to unveil the subtle hardware differences across various BSs, effectively capturing the unique BS energy fingerprint for more accurate predictions.
Figure 1: Illustration of “one-to-many” issue.
@@ IV Deep Learning Model Architecture and Training
In this section, we present our methodology, encompassing the input processing technique, the architecture of our deep learning model, and our innovative masked training approach, aimed at enhancing generalization across different BSs to address the one-to-many issues presented in Section III.
@@ IV-A Input Feature Processing
Our input feature processing is designed to address the diversity of input features through distinct feature encoding methods. Categorical inputs, except for BSID, are processed using One-Hot encoding, which converts them into a binary format, maintaining categorical distinctions. Numerical inputs, on the other hand, are inputted directly into the model in their original form to utilize their inherent magnitude variations. However, for numerical parameters like Antennas, Bandwidth, and Frequency with limited value ranges, we also apply One-Hot encoding to capture subtleties.
For the BSID feature, characterized by its extensive range of classes, an embedding technique is used, transforming the information into a compact numerical vector, thereby balancing dimensionality with information retention.
This tailored approach to input processing has led to improved performance, as detailed in Section IV. The encoding process adopted for all input parameters is outlined in Table I.
TABLE I: Input parameters.
@@ IV-B Deep Learning Model Design
Fig. 2(a) illustrates the architecture of our proposed framework. At its core, the model employs a multilayer perceptron (MLP) structure, and includes specialized modules for embedding and adaptive weight adjustment. The model consists of two hidden layers: the first layer contains 128 neurons, while the second layer comprises 64 neurons. The final layer directly outputs the predictions.
@@
(a) Deep learning model
(b) Training and inference phase
@@
To further enhance feature selection and to more effectively identify key features in the model, we propose an optimized network structure, incorporating a novel attention mechanism module, inspired by attention mechanisms utilized in Natural Language Processing (NLP) and Computer Vision (CV) applications. Our module, named adaptive re-weighting layer (ARL), enhances model predictions by assigning greater weights to the most significant input features. Our ARL module operates by initially compressing the dimensionality of the neural network input, $x$, using a linear layer. This compression is followed by a subsequent expansion in dimensionality. To facilitate non-linear processing, a rectified linear unit (ReLU) function is incorporated at the intermediate stage. The process completes with the application of a sigmoid activation function, yielding a vector that matches the input’s dimensionality. This vector comprises values in the range of 0 to 1, effectively redistributing weights across the input features. This mechanism is encapsulated in the following equation:
@@ IV-C Training with Random BSID Masking
To enhance our model’s generalization across unknown BSs, we have integrated a distinct BSID category for such BSs. As illustrated in Fig. 2(b), during the training phase, 30% of the samples in each epoch are randomly assigned this ‘unknown BS’ identifier. This technique is designed to enable the model to handle unknown BSIDs during the inference phase, enhancing its ability to generalize and avoid overfitting to BSs in the test set.
@@
where $y_{i}$ represents the i-th actual energy consumption sample, and $\widehat{y}_{i}$ denotes the model’s corresponding estimation.
Training was conducted over 1000 epochs with a batch size of 512. Additionally, an early stopping mechanism was adopted, halting training based on the validation set performance to further prevent overfitting.
@@ V-A Overall performance
Our model underwent extensive training, spanning over 1000 epochs, on an RTX 3090 GPU, with each epoch averaging a duration of one second. The trained model was tested to estimate the energy consumption of the 10,076 samples included in the testing set, where it achieved remarkable performance, evidenced by a MAPE value of 4.9811%.
@@ V-B BSID Generalization Capabilities
In this section, we discuss the strategies employed to enhance our model’s ability to generalize across different BSs by varying the processing of the BSID. We examined the impact of utilizing either One-Hot encoded BSID and embedding BSID or input features on model performance.
As indicated in Table II, employing One-Hot BSID in our model led to substantial predictive errors for unknown BSs. Without masked training, the model’s errors were more pronounced and performance worse than omitting BSID. However, using masked training significantly improved generalization for unknown BSs while maintaining accuracy for known BSs. Furthermore, the adoption of BSID embedding, in conjunction with masked training, led to a significant accuracy boost for unknown BSs, outperforming the results attained with One-Hot encoding. We also conducted experiments using the method proposed by [14] as a baseline. Due to its omission of the BSID as an input feature and a different encoding approach, it exhibited poor performance on this dataset.
To understand these dynamics, we employed Uniform Manifold Approximation and Projection (UMAP) for dimensionality reduction, compressing each BS’s embedding into a 2-dimensional space for visual analysis. As illustrated in Fig. 3, the 2-dimension visualization of BSID embeddings is color-coded according to RUType to demonstrate that it successfully captures the relationships between different BSs. This visualization reveals clustering patterns and proximity in the embeddings space between different RUTypes, showing that BSs of the same RUType are closer to each other in this space.
@@
Fig. 4(a) and 4(b) show the energy consumption predictions outputted by our model for a specific known RUType, comparing scenarios where BSID embedding is either included or omitted. When the BSID is excluded, the model converges towards a generalized pattern, aligning energy consumption estimates with the average behavior observed across all BSs for each RUType. In contrast, the integration of BSID embeddings enables the model to capture and incorporate the distinctive attributes of each individual BS thereby refining the precision of its energy consumption predictions.
(a) Without BSID
(b) With BSID embedding
(c) Different encoding methods for ABF
Figure 4: Comparative visualization of predictions for a known RUType dataset: (a) Model without BSID, (b) Model with embedding BSID, and (c) One-Hot vs. Numerical encoding for ABF on an unknown RUType dataset.
TABLE II: Comparison of BSID encoding methods, assessing the impact of random masking on model parameters and MAPE (%). Best performances are highlighted in bold. RM is used to denote Random Masking.
Methods (Dim) Parameters Unknown-BS Known-BS All-BS Embedding BSID w/ RM(204) (Ours) 105209 6.6863 4.0 4.9811 Embedding BSID w/o RM(204) 105209 21.8093 3.9984 11.4294 One-Hot BSID w/ RM(1064) 286219 9.44 3.9452 6.1313 One-Hot BSID w/o RM(1064) 286219 20.9739 4.0732 11.2645 No BSID(140) 28757 8.1589 9.2191 8.7538 Baseline[14](55) 2871 13.9083 11.3856 12.7587
@@ V-C Performance Comparison of One-Hot Encoding Combinations
In this section, we analyze the performance of our model in relation to different methods for incorporating some of the numerical features. Despite the inherent numerical nature of frequency, bandwidth, and the number of antennas, a consistent and clear linear relationship between these features and energy consumption is not always evident across all BSs, which may influence the precision of our model.
To address this issue, we explored treating these features both as numerical and as categorical variables. This approach entailed conducting a series of comparative analyses, varying the feature configurations between purely numerical and categorically encoded using One-Hot encoding. The primary aim of our experiments was to identify the configuration that achieves the lowest MAPE. The results of these investigations are detailed in Table III. In this table, “A”, “B”, and “F” denote Antennas, Bandwidth, and Frequency respectively. The configurations such as “ABF”, “BF”, “AF”, “AB”, “A”, “B”, and “F” suggest that these features were treated as categorical and preprocessed using One-Hot encoding. Notably, the “ABF” configuration emerges as the most effective, achieving the lowest MAPE for both known and unknown BSs.
Furthermore, Fig. 4(c) shows a visualization of energy consumption predictions, comparing the outcomes when the number of antennas, bandwidth, and frequency are processed as numerical features versus One-Hot features for a new RUType. This comparison illustrates that treating these inputs as categorical via One-Hot encoding yields energy consumption estimates that more accurately align with the groundtruth.
TABLE III: Performance comparison of different One-Hot encoding combinations on MAPE (%).
Combinations (Dim) Unknown-BS Known-BS All-BS
ABF (204) (Ours) 6.6863 4.0 4.9811
AB (172) 15.3823 4.5665 9.4701
AF (188) 13.0811 4.0847 7.9097
BF (199) 11.7292 4.1652 7.34
A (156) 20.094 4.7138 11.3221
B (167) 12.7202 4.2781 7.3687
F (183) 11.1229 4.3239 7.2736
Numerical (151) 13.7479 4.0733 8.3435

### v2

@@ Abstract
The introduction of 5G technology has revolutionized communications, enabling unprecedented capacity, connectivity, and ultra-fast, reliable communications. However, this leap has led to a substantial increase in energy consumption, presenting a critical challenge for network sustainability. Accurate energy consumption modeling is essential for developing energy-efficient strategies, enabling operators to optimize resource utilization while maintaining network performance. To address this, we propose a novel deep learning model for 5G base station energy consumption estimation based on a real-world dataset. Unlike existing methods, our approach integrates the Base Station Identifier (BSID) as an input feature through an embedding layer, capturing unique energy patterns across different base stations. We further introduce a masked training method and an attention mechanism to enhance generalization and accuracy. Experimental results show significant improvements, reducing Mean Absolute Percentage Error (MAPE) from 12.75% to 4.98%, achieving over 60% performance gain compared to existing models. The source code for our model is available at GitHub Repository.
@@ I Introduction
The fifth-generation (5G) network has revolutionized communication technologies, enabling transformative applications such as virtual reality, vehicular networks, and the Internet of Things (IoT). However, this technological advancement has also led to a significant increase in energy consumption. Compared to the fourth-generation (4G) network, 5G networks consume approximately three times more energy [1]. For mobile network operators, energy costs now represent around 22.5% of total operational expenses [2], and this proportion continues to grow. A major contributor to this energy consumption is the radio access network (RAN), with base stations (BSs) alone accounting for over 70% of the total network energy usage [3].
Although base stations (BSs) are inherently energy-intensive, their energy consumption can be optimized by dynamically disabling certain hardware components based on traffic load. Accurate energy consumption modeling is crucial for implementing such energy-saving strategies. Previous studies have extensively investigated the impact of BS architectures and configurations on power consumption. For instance, [4] proposed an analytical model demonstrating a linear relationship between BS output and power consumption. This model was later extended in [5] to incorporate massive multiple-input multiple-output (mMIMO) architectures and energy-efficient techniques. Further refinements were made in [6], which accounted for the linear increase in power consumption with additional mMIMO transceivers. A comprehensive framework was also introduced in [7], considering factors such as mMIMO architecture, downlink and uplink phases, and the number of multiplexed users. Additionally, the 3GPP technical report TR 38.864 [8] provides a standardized energy consumption model for 5G networks, focusing on energy-saving techniques such as dynamic resource allocation and power amplifier efficiency optimization. However, these models fail to fully capture the complexity of real-world 5G radio units, particularly the heterogeneity of hardware and environmental factors.
@@
Most recently, SRCON (Simulated Reality of Communication Networks) has emerged as a revolutionary paradigm in network simulation. Detailed in  [9, 10], SRCON leverages an innovative mix of white-box and black-box models to mimic the stochastic behaviors observed in real-world 4G/5G mobile networks, offering a robust framework for accurate simulation. Within this framework, [11, 12] develops a machine learning model, trained on extensive network measurement data, which significantly improves the accuracy of energy consumption estimations compared to traditional mathematical modeling.
However, these approaches typically overlook the domain shifts caused by variations in hardware configurations and physical environments. For instance, different base stations may have completely distinct equipments and physical environments, yet such differences are not adequately represented in the input features of current models. This domain discrepancy inevitably leads to prediction biases. The observed phenomenon - where data from the same base station shows consistent patterns while significant variations exist across different stations.
To address these limitations, we propose a data-driven approach based on a real-world 5G radio unit dataset from the ITU 5G Base Station Energy Consumption Modelling Challenge [13]. The dataset exhibits a characteristic domain shift problem, where identical input features correspond to varying energy consumption patterns across different base stations. Unlike existing methods, our approach incorporates the Base Station Identifier (BSID) as a key feature to capture the domain-specific characteristics. We utilize an embedding layer to represent the BSID, which provides a more compact representation compared to sparse one-hot encoding. To further improve robustness, we introduce a masked training technique to handle unseen BSIDs in test data. Additionally, an attention mechanism is employed to automatically select the most relevant features for energy consumption estimation. Experimental results demonstrate that our approach reduces the Mean Absolute Percentage Error (MAPE) from 12.75% to 4.98%, achieving a performance improvement of over 60% compared to state-of-the-art methods [12].
@@
In this section, we introduce the 5G Radio Unit Dataset used in our study. The dataset consists of 102,705 hourly measurements collected over a period of 8 days. Each sample includes a variety of features, as detailed in Table I, which are classified into three main categories along with energy consumption:
@@
The dataset is split into training and testing sets to test our model’s generalization capabilities, with certain BSs only present in the testing set. The training set consists of 92,629 samples. The testing set includes 10,076 samples, of which 3,067 are from BSs not encountered during training (cross-domain), and 7,009 are from BSs that appear in the training set (in-domain).
@@ III Problem Statement
By analysis of the collected data, there exists a “one-to-many” issue in the datasets, meaning that different energy consumption values are reported for identical input feature values. This issue arises from two primary origins: i) the absence of certain unknown features, leading to a loss of information; ii) errors that occur during the measurement of energy consumption. Fig. 1 is a scatter plot showing the energy consumption values of selected data samples plotted against the load of the primary cell, which is identified as one of the most important features in [12]. Three key observations can be made from Fig. 1 as listed below:
  • Energy consumption exhibits a positive correlation with the load of the primary cell, where higher loads tend to correspond to higher energy consumption.
  • Without distinguishing BSID, as depicted by the gray area in the figure, a common load value corresponds to a high variance in energy consumption.
@@
Combining these observations, we conclude that distinguishing BSID enhances our ability to identify patterns in energy consumption variations across different BSs. This gain comes from the fact that incorporating BSID as a model input feature enables us to unveil the subtle hardware differences across various BSs, effectively capturing the unique BS energy fingerprint for more accurate predictions.
Figure 1: Illustration of the “one-to-many” issue. Different colors represent sample points from different BS while sharing the same input features.
To mathematically model the relationship between input features and energy consumption, we define a mapping function $f$ where $y\in\mathbb{R}$ denotes the energy consumption, and $\boldsymbol{x}\in\mathbb{R}^{n}$ denotes the input features, formed by concatenating various features and BSID, as shown in Table I. Since deriving $f(\boldsymbol{x})$ analytically is challenging, we propose using a neural network $f(\boldsymbol{x};\theta)$ with trainable parameters $\theta$:
$y=f(\boldsymbol{x};\theta).$ (1)
By including BSID in $\boldsymbol{x}$, the model captures unique base station characteristics, reducing uncertainty from unknown factors.
In practical deployment scenarios, the proposed method allows for the use of alternative features during local training, rather than being limited to the features used in the dataset. The requirements for the amount of training samples can be referenced in [12].
@@ IV Deep Learning Model Architecture and Training
In this section, we present our methodology, which includes input processing techniques, the architecture of our deep learning model, and a novel masked training approach designed to improve generalization across different BSs. This methodology addresses the one-to-many issues outlined in Section III.
@@ IV-A Input Feature Processing
Our input feature processing employs distinct encoding methods to handle the diversity of input features effectively. Categorical inputs (excluding BSID) are processed using One-Hot encoding to preserve categorical distinctions. Numerical inputs are used directly to leverage their magnitude variations, except for features like Antennas, Bandwidth, and Frequency, which have limited value ranges and are One-Hot encoded to capture subtle variations. For BSID, which has a large number of classes, we use embedding techniques to transform it into a compact numerical vector, balancing dimensionality and information retention. This tailored approach has significantly improved model performance, as discussed in Section IV. The encoding methods for all input features are summarized in Table I.
TABLE I: Input Features.
@@ IV-B Deep Learning Model Design
Fig. 2(left) illustrates the architecture of our proposed framework. At its core, the model employs a multilayer perceptron (MLP) structure, and includes specialized modules for embedding and adaptive weight adjustment. The model consists of two hidden layers: the first layer contains 128 neurons, while the second layer comprises 64 neurons. The final layer directly outputs the predictions.
@@
To enhance feature selection and identify key elements within the feature vector, we propose an optimized network structure with a novel attention mechanism, the adaptive re-weighting layer (ARL). ARL improves predictions by assigning different weights to each feature element. The module first compresses the input, $x$ via a linear layer, then expands it back to the original dimensionality. A rectified linear unit (ReLU) function is applied in the intermediate stage, followed by a sigmoid activation, producing a vector of values between 0 and 1 to re-weight the input features. This process is described by the following equation:
@@ IV-C Training with Random BSID Masking
To enhance our model’s generalization across unknown BSs, we have integrated a distinct BSID category for such BSs. As illustrated in Fig. 2(right), during the training phase, 30% of the samples in each epoch are randomly assigned this ‘unknown BS’ identifier. This technique is designed to enable the model to handle unknown BSIDs during the inference phase, enhancing its ability to generalize and avoid overfitting to BSs in the test set.
@@
where $y_{i}$ represents the i-th measured energy consumption sample, and $\widehat{y}_{i}$ denotes the model’s corresponding estimation.
The model was trained for 1000 epochs with a batch size of 512. After each epoch, the model was evaluated on the test set, and the epoch with the best test performance was selected as the final model. This approach ensures that the model’s performance is reported under its optimal conditions.
In practical eployment scenarios, the masked training strategy can be employed for datasets with different input features. During training, the model alternates between using and omitting specific feature values across epochs, enabling adaptation to unseen values while maintaining performance on observed values. This enhances generalization to new or incomplete data during inference.
@@ V-A Overall performance
Our model was trained for over 1000 epochs. The trained model was tested on 10,076 samples from the testing set and achieved remarkable performance, with a MAPE value of 4.9811%.
@@ V-B BSID Generalization Capabilities
In this section, we discuss strategies to enhance our model’s generalization across different BSs by varying the processing of BSID. We compare two approaches: using One-Hot encoded BSID versus embedded BSID alongside other input features.
As shown in Table II, using One-Hot encoded BSID resulted in significant predictive errors for unknown BSs. Without masked training, the model performed worse than omitting BSID entirely. However, masked training improved generalization for unknown BSs while maintaining accuracy for known BSs. Further, combining BSID embedding with masked training significantly boosted accuracy for unknown BSs, outperforming One-Hot encoding. For comparison, we also tested the method proposed by [11], which omits BSID and uses a different encoding approach, but it performed poorly on this dataset.
To analyze these results, we used UMAP for dimensionality reduction, compressing BS embeddings into 2D for visualization. Fig. 3 shows that BSs of the same RUType cluster together in the embedding space, indicating that the embedding effectively captures structural similarities.
@@
Fig. 4(a) compares energy consumption predictions for a known RUType with and without BSID embedding. When BSID is omitted, the model converges to a generalized pattern, predicting energy consumption based on the average behavior of all BSs for that RUType. In contrast, incorporating BSID embedding allows the model to capture the unique characteristics of individual BSs, significantly improving prediction accuracy.
(a) Without BSID (left) and with BSID embedding (right)
(b) Different encoding methods for ABF
Figure 4: Comparative visualization of predictions for a known RUType dataset: (a) Model without BSID, model with embedding BSID, and (b) One-Hot vs. Numerical encoding for ABF on an unknown RUType dataset.
@@ V-C Performance Comparison of One-Hot Encoding Combinations
In this section, we evaluate our model’s performance by exploring different methods for incorporating numerical features such as frequency, bandwidth, and the number of antennas. Although these features are inherently numerical, their values in the dataset are sparse (i.e., they have limited and non-dense distributions). This sparsity makes it challenging to establish a clear linear relationship between these features and energy consumption across all BSs, potentially impacting model accuracy.
Given the sparse nature of these features, we investigated whether treating them as categorical variables, rather than numerical, could improve performance. Specifically, we compared their treatment as numerical features versus categorical features using One-Hot encoding. The goal was to identify the configuration that achieves the lowest MAPE. Table II summarizes the results, where “A”, “B”, and “F” represent Antennas, Bandwidth, and Frequency, respectively. Configurations such as “ABF”, “BF”, “AF”, “AB”, “A”, “B”, and “F” indicate categorical treatment with One-Hot encoding. The “ABF” configuration emerged as the most effective, achieving the lowest MAPE for both known and unknown BSs. This outcome underscores the advantage of using One-Hot encoding for sparse numerical features.
Furthermore, Fig. 4(b) visualizes energy consumption predictions for a new RUType, comparing numerical and One-Hot encoding. The results demonstrate that categorical encoding aligns more closely with groundtruth data, particularly for sparse features like ABF. This confirms that the sparsity of these numerical features in the dataset is a key factor driving the superior performance of One-Hot encoding in this context.